OpenStudy (anonymous):

Which of the following equations are identities? A) sin^2 x - cos^2 x = 1 B) tan^2 x - sec^2 x -1 C) sin^2 x = 1 - cos^2 x D) cot^2 x + csc^2 x = 1

OpenStudy (anonymous):

Will medal!

OpenStudy (johnweldon1993):

Always remember \[\large sin^2(x) + cos^2(x) = 1\] you can derive any other identities from that

OpenStudy (anonymous):

\[\sin ^2x+\cos ^2 x=1 ......(1)\] rearrange and find the solution. divide each term of (1) by \[\sin ^2x\] and rearrange again divide each term of (1) by \[\cos^2 x\] and rearrange

Nnesha (nnesha):

divide..?